# EEG-Deformer: A Dense Convolutional Transformer for Brain-computer Interfaces

### v1

@@ Abstract
Effectively learning the temporal dynamics in electroencephalogram (EEG) signals is challenging yet essential for decoding brain activities using brain-computer interfaces (BCIs). Although Transformers are popular for their long-term sequential learning ability in the BCI field, most methods combining Transformers with convolutional neural networks (CNNs) fail to capture the coarse-to-fine temporal dynamics of EEG signals. To overcome this limitation, we introduce EEG-Deformer, which incorporates two main novel components into a CNN-Transformer: (1) a Hierarchical Coarse-to-Fine Transformer (HCT) block that integrates a Fine-grained Temporal Learning (FTL) branch into Transformers, effectively discerning coarse-to-fine temporal patterns; and (2) a Dense Information Purification (DIP) module, which utilizes multi-level, purified temporal information to enhance decoding accuracy. Comprehensive experiments on three representative cognitive tasks consistently verify the generalizability of our proposed EEG-Deformer, demonstrating that it either outperforms existing state-of-the-art methods or is comparable to them. Visualization results show that EEG-Deformer learns from neurophysiologically meaningful brain regions for the corresponding cognitive tasks. The source code can be found at https://github.com/yi-ding-cs/EEG-Deformer.
Index Terms:
@@
I Introduction
Brain-computer interface (BCI) technology facilitates direct communication between the brain and machines using electroencephalography (EEG) [1]. A standard BCI system usually consists of four key components: data acquisition, pre-processing, classification, and feedback [2]. BCIs are employed in various practical applications, such as stroke rehabilitation [3] and emotion regulation in mental health treatments [4].
@@ I Introduction
Recently, numerous deep learning-based approaches have been developed for decoding brain activities from EEG signals. These approaches can be broadly categorized into two types: methods using hand-crafted features and methods using EEG directly as input. The former [8, 9, 10, 11] involve extracting various types of features from EEG signals for neural network input. Meanwhile, CNN-based methods, which leverage automatic feature learning, typically process EEG as 2-D time series [12, 13, 14]. Given that EEG data is inherently graph-structured, with nodes representing electrodes and connections based on spatial distance, functional connectivity, or learned relationships, graph-based methods have gained popularity [8, 15]. However, using features as input may result in the loss of fine-grained temporal information by averaging data along the temporal dimensions. Additionally, CNN-based methods may fail to capture long-term temporal dependencies by employing CNN kernel along temporal dimension.
@@
Furthermore, to efficiently utilize multi-level temporal information from intermediate HCT layers, we have designed a Dense Information Purification (DIP) module in our EEG-Deformer. This module enables the dense transmission of multi-level representations from HCT layers to the final representation. Differing from the skip connections in DenseNet [20], we introduce an Information Purification Unit (IP-Unit) that transforms the fine-grained representations from FTL branches using a logarithmic (log) power operation [15] and elegantly fuses these latent representations into the final representation. Because log power can represent the amount of activity in filtered signals and reduce dimensions [21], our IP-Unit not only retains critical frequency characteristics related to brain activity but also effectively reduces the number of learnable parameters. As illustrated in Figure 1, a series of IP-Units have been added to the intermediate layers. These units progressively encode discriminative temporal representations. We will validate these enhancements in Sec. V-B to V-E.
@@
Figure 2: The network structure of EEG-Deformer. EEG-Deformer consists of three main parts: (1) Shallow feature encoder, (2) Hierarchical coarse-to-fine-Transformer (HCT), and (3) Dense information purification (DIP). The fine-grained representations from each HCT will be passed to Information Purification Unit (IP-Unit) and concatenated to the final embedding.
II Related Work
II-A CNNs for EEG Decoding
CNNs have emerged as potent tools in BCI applications, adept at directly learning from EEG data. Schirrmeister et al. [12] introduced DeepConvNet, which employs a dual-stage spatial and temporal convolution layer for EEG data feature extraction and classification. In parallel, Lawhern et al. [13] created EEGNet, utilizing depth-wise convolution with a kernel size of $(n,1)$, where $n$ signifies the number of EEG channels, to capture spatial information. Advancing this concept, TSception [14] leverages multi-scale convolutional kernels to decode EEG’s temporal dynamics and asymmetric spatial activations. However, the relatively short length of these 1-D CNN kernels in the temporal dimension slightly curtails their efficacy in learning long-term temporal patterns.
II-B Transformers for BCI
Transformers, known for their ability to learn long-term correlations in sequential data, have garnered increasing attention from researchers [16, 17, 18, 6]. Lee et al. [18] demonstrate that incorporating a self-attention module from Transformer architecture into an EEGNet-based CNN enhances the classification of imagined speech using EEG. EEG-Conformer [6], a compact convolutional Transformer, effectively merges local and global features for EEG decoding in a cohesive framework. However, most existing convolutional Transformers, which typically use CNNs for shallow feature extraction followed by Transformer blocks, may not efficiently learn the coarse- and fine-grained temporal information encoded in EEG signals. Additionally, they often overlook the multi-level temporal information available from different layers.
III Method
@@
III-A Shallow Feature Encoders
@@
Let’s denote an EEG sample as $\textit{{X}}\in\mathbb{R}^{c\times l}$, where $c$ is the number of EEG channels, and $l$ represents the number of data points along the time dimension. Inspired by neurophysiological knowledge that the brain has microstates lasting approximately 100 ms, the temporal CNN kernels are sized at $(1,0.1\times f_{s})$, where $f_{s}$ denotes the EEG’s sampling rate. For capturing spatial information from EEG, a spatial CNN kernel of size $(c,1)$ is utilized as suggested in [13], where $c$ is the number of EEG channels. Weight normalization is included following the approach in [22]. The number of CNN kernels is denoted by $k$. After activation by the ELU function, the learned features are max-pooled every two data points without overlapping. The tokenizer includes rearrange operation and a learnable position encoding. The size of the features is then rearranged into $k\times 0.5l$ to serve as tokens for the Transformers, which learn coarse-grained temporal information. Subsequently, a learnable position encoding, $\textit{{P}}\in\mathbb{R}^{k\times 0.5l}$, is added to the tokens [5]. Therefore, the encoded tokens can be represented as $\textit{{F}}\in\mathbb{R}^{k\times 0.5l}$.
@@
III-B Hierarchical Coarse-to-fine Transformer
@@
Let us suppose $\textit{{F}}_{i}\in\mathbb{R}^{k\times l_{t}^{i}}$ the input to the $i$-th HCT block. To capture the coarse-grained temporal dynamics of EEG signals, we treat the output of each CNN kernel as one token. By doing so, the long-term temporal information is explicitly included when we project the tokens into $\textit{{Q}}_{i}\in\mathbb{R}^{n_{head}\times k\times d_{attn}}$, $\textit{{K}}_{i}\in\mathbb{R}^{n_{head}\times k\times d_{attn}}$, and $\textit{{V}}_{i}\in\mathbb{R}^{n_{head}\times k\times d_{attn}}$ using the linear projection (LP) layer parameterized by $n_{head}$ set of $\textit{{W}}_{qkv}^{i}\in\mathbb{R}^{\frac{l_{t}^{i}}{2^{i-1}}\times 3d_{attn}}$. Before the LP layer, a max pooling layer is added to reduce the feature dimension of $\textit{{F}}_{i}$. The encoding process $\Phi(\cdot)$ can be formulized by
@@
We further propose a FTL module to capture the short-period temporal dynamics of EEG signals. To learn fine-grained temporal features, we opt for a 1D CNN, as its short kernel moves along the temporal dimension step by step, enabling the learning of short-period patterns. After a dropout layer, the learned representations are fed into a 1-D CNN layer, followed by a batch normalization layer, an ELU activation, and a max-pooling layer. The fine-grained temporal representations, denoted as $\bm{F}_{fg}^{i}$, can be calculated as follows:
@@
III-C Dense Information Purification
@@
TABLE I: Generalized cross-subject classification results from various methods on three benchmarking datasets. The best results are highlighted in bold and the next best are marked using underlines
Method Dataset I Dataset II Dataset III ACC std F1-macro std ACC std F1-macro std ACC std F1-macro std DGCNN 60.98 6.05 56.81 9.51 69.56 12.50 64.92 13.08 64.85 17.90 59.42 22.25 LGGNet 67.81 6.38 66.70 7.69 68.16 13.16 65.38 12.40 65.37 12.78 61.91 16.16 EEGNet 75.28 9.77 73.81 12.62 73.95 12.54 71.43 13.45 65.85 14.55 61.89 18.21 TSception 71.60 7.66 69.94 10.88 73.01 12.93 70.18 13.90 69.92 15.07 65.26 20.04 EEG-ViT 67.72 7.03 66.79 8.49 66.95 14.12 64.12 13.89 63.12 15.44 61.19 16.56 SSVEPformer 73.18 7.87 72.7 8.54 55.02 10.95 52.00 9.87 55.17 6.71 54.25 7.01 EEG-Transformer 75.39 7.92 74.93 8.61 65.53 11.2 63.86 12.84 64.46 14.91 64.28 14.94 EEG-Conformer 79.81 9.27 79.01 10.75 74.36 11.82 71.65 12.96 69.40 14.48 65.59 18.68 EEG-Deformer (ours) 82.72 8.00 82.36 8.52 79.32 7.87 75.83 11.54 73.18 15.63 69.99 19.80
TABLE II: Results of the ablation study conducted on three benchmarking datasets, with the best results marked in bold.
Method Dataset I Dataset II Dataset III ACC std F1-macro std ACC std F1-macro std ACC std F1-macro std w/o FTL 70.39 6.89 69.89 7.39 70.48 14.76 68.63 13.97 63.31 11.67 61.79 13.25 w/o Dense connection 75.84 7.24 75.40 7.67 70.85 14.11 68.86 14.22 72.46 14.73 69.30 18.88 w/o IP-Unit 72.01 7.26 71.28 8.06 73.28 12.10 70.98 12.64 67.29 11.75 64.48 14.73 EEG-Deformer (ours) 82.72 8.00 82.36 8.52 79.32 7.87 75.83 11.54 73.18 15.63 69.99 19.80
IV Experiment
IV-A Datasets
@@
Dataset I: Dataset I provides EEG signals recorded while subjects perform the Discrimination/Selection Response (DSR) task to assess cognitive attention. The experiment involves a total of 26 participants. Each subject undergoes three sessions, with each session consisting of multiple cycles of attention tasks, each lasting 40 seconds, and alternating with 20-second rest periods. Data collection includes recordings from 28 EEG channels, sampled at a rate of 1000 Hz. According to [15], the first 20 seconds of attention tasks are used to balance the classes. All three sessions are utilized in this study. Each EEG trial is segmented using a sliding window of 4 seconds with a 50% overlap.
Dataset II: Dataset II is designed to evaluate cognitive fatigue states in drivers through EEG signal analysis during a 90-minute driving task within a virtual reality (VR) driving environment. This dataset was compiled with the participation of 27 subjects. EEG data were captured using a 32-channel EEG system, with a sampling rate of 500 Hz. The official pre-processed data are used. Following [15], 11 subjects who had enough samples for each class are utilized. The data is down-sampled to 128 Hz. Each sample contains of 3-second EEG signals.
@@
Dataset III: Dataset III provides 19-channel EEG recordings from 36 subjects engaged in mental cognitive tasks, specifically performing serial subtraction, along with their corresponding baseline EEGs for reference. The official artifact-free data are used in this study. Each mental workload trial lasts for 60 seconds, and the last 60 seconds of the rest EEG are used as the low workload data. The trials are segmented using a 4-second sliding window with a moving step of 2 seconds.
IV-B Baselines

### v2

@@ Abstract
Effectively learning the temporal dynamics in electroencephalogram (EEG) signals is challenging yet essential for decoding brain activities using brain-computer interfaces (BCIs). Although Transformers are popular for their long-term sequential learning ability in the BCI field, most methods combining Transformers with convolutional neural networks (CNNs) fail to capture the coarse-to-fine temporal dynamics of EEG signals. To overcome this limitation, we introduce EEG-Deformer, which incorporates two main novel components into a CNN-Transformer: (1) a Hierarchical Coarse-to-Fine Transformer (HCT) block that integrates a Fine-grained Temporal Learning (FTL) branch into Transformers, effectively discerning coarse-to-fine temporal patterns; and (2) a Dense Information Purification (DIP) module, which utilizes multi-level, purified temporal information to enhance decoding accuracy. Comprehensive experiments on three representative cognitive tasks—cognitive attention, driving fatigue, and mental workload detection—consistently confirm the generalizability of our proposed EEG-Deformer, demonstrating that it either outperforms or performs comparably to existing state-of-the-art methods. Visualization results show that EEG-Deformer learns from neurophysiologically meaningful brain regions for the corresponding cognitive tasks. The source code can be found at https://github.com/yi-ding-cs/EEG-Deformer.
{IEEEkeywords}
@@
1 Introduction
[…]
Brain-computer interface (BCI) technology facilitates direct communication between the brain and machines using electroencephalography (EEG) [1]. A standard BCI system usually consists of four key components: data acquisition, pre-processing, classification, and feedback [2]. BCIs are employed in various practical applications, such as stroke rehabilitation [3], sleep stage detection[4, 5], and emotion regulation in mental health treatments [6].
@@ 1 Introduction
Recently, numerous deep learning-based approaches have been developed for decoding brain activities from EEG signals. These approaches can be broadly categorized into two types: methods using hand-crafted features and methods using EEG directly as input. The former [10, 11, 12, 13, 14] involve extracting various types of features from EEG signals for neural network input. Meanwhile, CNN-based methods, which leverage automatic feature learning, typically process EEG as 2-D time series [15, 16, 17]. Given that EEG data is inherently graph-structured, with nodes representing electrodes and connections based on spatial distance, functional connectivity, or learned relationships, graph-based methods have gained popularity [10, 18, 19]. However, using features as input may result in the loss of fine-grained temporal information by averaging data along the temporal dimensions. Additionally, CNN-based methods may fail to capture long-term temporal dependencies by employing CNN kernel along temporal dimension.
@@
Furthermore, to efficiently utilize multi-level temporal information from intermediate HCT layers, we have designed a Dense Information Purification (DIP) module in our EEG-Deformer. This module enables the dense transmission of multi-level representations from HCT layers to the final representation. Differing from the skip connections in DenseNet [24], we introduce an Information Purification Unit (IP-Unit) that transforms the fine-grained representations from FTL branches using a logarithmic (log) power operation [18] and elegantly fuses these latent representations into the final representation. Because log power can represent the amount of activity in filtered signals and reduce dimensions [25], our IP-Unit not only retains critical frequency characteristics related to brain activity but also effectively reduces the number of learnable parameters. As illustrated in Figure 1, a series of IP-Units have been added to the intermediate layers. These units progressively encode discriminative temporal representations. We will validate these enhancements in Sec. 5.2 to 5.5.
@@
Figure 2: The network structure of EEG-Deformer. EEG-Deformer consists of three main parts: (1) Shallow feature encoder, (2) Hierarchical coarse-to-fine-Transformer (HCT), and (3) Dense information purification (DIP). The fine-grained representations from each HCT will be passed to Information Purification Unit (IP-Unit) and concatenated (C) to the final embedding.
2 Related Work
2.1 CNNs for EEG Decoding
Convolutional Neural Networks (CNNs) have become powerful tools in BCI applications, excelling in learning directly from EEG data [15, 16, 17, 26]. Schirrmeister et al. [15] introduced DeepConvNet, which incorporates a two-stage spatial and temporal convolution layer to facilitate EEG feature extraction and classification. Similarly, Lawhern et al. [16] developed EEGNet, using depth-wise convolution with a kernel size of $(n,1)$—where $n$ represents the number of EEG channels—to capture spatial features. Building on these methods, TSception [17] employs multi-scale convolutional kernels to decode temporal dynamics and asymmetric spatial activations within EEG signals. However, the relatively short length of these 1-D CNN kernels in the temporal dimension limits their ability to capture long-term temporal patterns effectively.
2.2 Transformers for BCI
Transformers, renowned for capturing long-term dependencies in sequential data, have gained considerable attention in EEG research [20, 21, 22, 8, 27, 28]. Lee et al. [22] showed that integrating a self-attention module from the Transformer architecture into an EEGNet-based CNN improves classification accuracy for imagined speech from EEG data. EEG-Conformer [8], a streamlined convolutional Transformer model, effectively combines local and global features for EEG decoding within a unified structure. However, most existing convolutional Transformers, which typically leverage CNNs for shallow feature extraction followed by Transformer blocks, may struggle to learn both coarse- and fine-grained temporal patterns in EEG signals. Furthermore, they frequently miss out on capturing multi-level temporal information available across different layers.
3 Method
@@
3.1 Shallow Feature Encoders
@@
Let’s denote an EEG sample as $\textit{{X}}\in\mathbb{R}^{c\times l}$, where $c$ is the number of EEG channels, and $l$ represents the number of data points along the time dimension. Inspired by neurophysiological knowledge that the brain has microstates lasting approximately 100 ms, the temporal CNN kernels are sized at $(1,0.1\times f_{s})$, where $f_{s}$ denotes the EEG’s sampling rate. For capturing spatial information from EEG, a spatial CNN kernel of size $(c,1)$ is utilized as suggested in [16], where $c$ is the number of EEG channels. Weight normalization is included following the approach in [29]. The number of CNN kernels is denoted by $k$. After activation by the ELU function, the learned features are max-pooled every two data points without overlapping. The tokenizer includes rearrange operation and a learnable position encoding. The size of the features is then rearranged into $k\times 0.5l$ to serve as tokens for the Transformers, which learn coarse-grained temporal information. Subsequently, a learnable position encoding, $\textit{{P}}\in\mathbb{R}^{k\times 0.5l}$, is added to the tokens [7]. Therefore, the encoded tokens can be represented as $\textit{{F}}\in\mathbb{R}^{k\times 0.5l}$. This step can be formularized as
@@
3.2 Hierarchical Coarse-to-fine Transformer
@@
Let us suppose $\textit{{F}}_{i}\in\mathbb{R}^{k\times l_{t}^{i}}$ the input to the $i$-th HCT block. To capture the coarse-grained temporal dynamics of EEG signals, we treat the output of each CNN kernel as one token. By doing so, the long-term temporal information is explicitly included when we project the tokens into $\textit{{Q}}_{i}\in\mathbb{R}^{n_{head}\times k\times d_{attn}}$, $\textit{{K}}_{i}\in\mathbb{R}^{n_{head}\times k\times d_{attn}}$, and $\textit{{V}}_{i}\in\mathbb{R}^{n_{head}\times k\times d_{attn}}$ using the linear projection (LP) layer parameterized by $n_{head}$ set of $\textit{{W}}_{qkv}^{i}\in\mathbb{R}^{\frac{l_{t}^{i}}{2^{i-1}}\times 3d_{attn}}$. Before the LP layer, a max pooling layer with a pooling size of 2 and a step of 2 is added to reduce the feature dimension of $\textit{{F}}_{i}$. The encoding process $\Phi(\cdot)$ can be formularized by
@@
We further propose a FTL module to capture the short-period temporal dynamics of EEG signals. To learn fine-grained temporal features, we opt for a 1-D CNN, as its short kernel moves along the temporal dimension step by step, enabling the learning of short-period patterns. After a dropout layer, the learned representations are fed into a 1-D CNN layer, followed by a batch normalization layer, an ELU activation, and a max-pooling layer. The fine-grained temporal representations, denoted as $\bm{F}_{fg}^{i}$, can be calculated as follows:
@@
3.3 Dense Information Purification
@@
Table 1: Generalized cross-subject classification results from various methods on three benchmarking datasets. The best results are highlighted in bold and the next best are marked using underlines
Method Dataset I: Attention Dataset II: Fatigue Dataset III: Mental Workload ACC std F1-macro std ACC std F1-macro std ACC std F1-macro std DGCNN 60.98 6.05 56.81 9.51 69.56 12.50 64.92 13.08 64.85 17.90 59.42 22.25 LGGNet 67.81 6.38 66.70 7.69 68.16 13.16 65.38 12.40 65.37 12.78 61.91 16.16 EEGNet 75.28 9.77 73.81 12.62 73.95 12.54 71.43 13.45 65.85 14.55 61.89 18.21 TSception 71.60 7.66 69.94 10.88 73.01 12.93 70.18 13.90 69.92 15.07 65.26 20.04 EEG-ViT 67.72 7.03 66.79 8.49 66.95 14.12 64.12 13.89 63.12 15.44 61.19 16.56 SSVEPformer 73.18 7.87 72.7 8.54 55.02 10.95 52.00 9.87 55.17 6.71 54.25 7.01 EEG-Transformer 75.39 7.92 74.93 8.61 65.53 11.2 63.86 12.84 64.46 14.91 64.28 14.94 EEG-Conformer 79.81 9.27 79.01 10.75 74.36 11.82 71.65 12.96 69.40 14.48 65.59 18.68 EEG-Deformer (ours) 82.72 8.00 82.36 8.52 79.32 7.87 75.83 11.54 73.18 15.63 69.99 19.80
Table 2: Results of the ablation study conducted on three benchmarking datasets, with the best results marked in bold.
Method Dataset I: Attention Dataset II: Fatigue Dataset III: Mental Workload ACC std F1-macro std ACC std F1-macro std ACC std F1-macro std w/o FTL 70.39 6.89 69.89 7.39 70.48 14.76 68.63 13.97 63.31 11.67 61.79 13.25 w/o Dense connection 75.84 7.24 75.40 7.67 70.85 14.11 68.86 14.22 72.46 14.73 69.30 18.88 w/o IP-Unit 72.01 7.26 71.28 8.06 73.28 12.10 70.98 12.64 67.29 11.75 64.48 14.73 EEG-Deformer (ours) 82.72 8.00 82.36 8.52 79.32 7.87 75.83 11.54 73.18 15.63 69.99 19.80
4 Experiment
4.1 Datasets and Pre-processing
@@
Dataset I: Dataset I provides EEG signals recorded while subjects perform the Discrimination/Selection Response (DSR) task to assess cognitive attention. The experiment involves a total of 26 participants. Each subject undergoes three sessions, with each session consisting of multiple cycles of attention tasks, each lasting 40 seconds, and alternating with rest periods of 20 seconds. Data collection includes recordings from 28 EEG channels, sampled at a rate of 1000 Hz. BrainAmp EEG amplifier (Brain Products GmbH, Gilching, Germany) was used to collect the EEG signals. For the pre-processing, a band-pass filter ranging from 0.5 to 50 Hz was applied to eliminate low and high-frequency noise, following the methodology in [11]. Eye movement artifacts were removed using the automatic Independent Component Analysis (ICA) Electrooculography (EOG) removal method in the MNE toolbox [35]. The data were then downsampled to 200 Hz. In accordance with [18], only the first half of each attention trial (20 seconds) was used to balance the samples between attention and inattention (rest). Each trial was further segmented into 4-second segments with a 50% overlap. All three sessions are utilized in this study.
Dataset II: Dataset II is designed to evaluate cognitive fatigue states in drivers through EEG signal analysis during a 90-minute driving task within a virtual reality (VR) driving environment. This dataset was compiled with the participation of 27 subjects. EEG data were captured using a 32-channel Scan SynAmps2 Express system (Compumedics Ltd., VIC, Australia), with a sampling rate of 500 Hz. The official pre-processed data are used. For pre-processing, the raw EEG signals were filtered by a band-pass filter from 1 to 50 Hz. Eye blinks were manually removed through visual inspection. Ocular and muscular artifacts were removed using the Automatic Artifact Removal (AAR) method in EEGLab [36]. Following [18], the data were downsampled to 128 Hz. For calculating fatigue levels, we also adhered to the approach outlined in [18]. EEG trials were defined as the 3 seconds of EEG data preceding the onset of lane-departure events. Reaction time (RT) was used to measure fatigue levels for these EEG trials. RT was defined as the interval from the onset of the lane-departure event to the onset of the counter-steering event. The local RT ($RT_{l}$) of a trial was defined as the RT for that specific trial. The global RT ($RT_{g}$) of a trial was the mean of the local RTs of all trials within a 90-second window before the current trial. The 5th percentile of all local RTs in the entire session was selected as the alert RT ($RT_{a}$). The labeling process was defined as follows:
$y=\left\{\begin{matrix}0&RT_{l}>2.5*RT_{a}\&\&RT_{g}>2.5*RT_{a}\\ 1&RT_{l}<1.5*RT_{a}\&\&RT_{g}<1.5*RT_{a}\end{matrix}\right.$ (10)
where 0 represent the fatigue class and 1 represent the non-fatigue class. Following [18], 11 subjects who had enough samples for each class are utilized.
@@
Dataset III: Dataset III provides 19-channel EEG recordings from 36 subjects engaged in mental cognitive tasks, specifically performing serial subtraction, along with their corresponding baseline EEG for reference. The official artifact-free data are used in this study. The EEG were recorded using a Neurocom EEG 23-channel system (Ukraine, XAI-MEDICA). Electrodes were positioned on the scalp according to the International 10-20 system, and all electrodes were referenced to ear reference electrodes. A high-pass filter with a 30 Hz cut-off frequency and a power line notch filter (50 Hz) were applied. ICA was used to eliminate artifacts. Each mental workload trial lasts for 60 seconds, and the last 60 seconds of the rest EEG are used as the low workload data. The trials are segmented using a 4-second sliding window with a moving step of 2 seconds.
Table 3: Effects of different information purification methods on three benchmarking datasets, with the best results highlighted in bold.
Method Dataset I: Attention Dataset II: Fatigue Dataset III: Mental Workload ACC std F1-macro std ACC std F1-macro std ACC std F1-macro std Using mean 77.35 6.99 76.99 7.33 73.82 12.13 71.72 12.93 68.82 15.44 65.47 18.90 Using std 77.54 7.54 77.04 8.07 73.94 14.42 72.34 14.77 72.27 15.20 69.19 19.00 Using power 82.72 8.00 82.36 8.52 79.32 7.87 75.83 11.54 73.18 15.63 69.99 19.80
Table 4: Effects of various information purification locations on three benchmarking datasets, with the best results highlighted in bold.
Method Dataset I: Attention Dataset II: Fatigue Dataset III: Mental Workload ACC std F1-macro std ACC std F1-macro std ACC std F1-macro std At $\bm{F}_{i+1}$ 80.28 7.46 79.81 8.03 77.25 7.14 73.98 9.96 69.16 15.46 66.35 18.19 At $\bm{F}_{cg}^{i}$ 72.90 7.30 72.05 8.35 44.72 16.48 29.96 8.30 71.41 15.30 68.56 18.50 At $\bm{F}_{fg}^{i}$ 82.72 8.00 82.36 8.52 79.32 7.87 75.83 11.54 73.18 15.63 69.99 19.80
4.2 Baselines